A rectangular garden is three times as long as it is wide. if the area of the garden is 1323 ft^2, find the length and width of the garden.
Answer by nerdybill(7384)

A rectangular garden is three times as long as it is wide. if the area of the garden is 1323 ft^2, find the length and width of the garden.
.
Let w = width
then
3w = length
.
w(3w) = 1323
3w^2 = 1323
w^2 = 441
w = 21 feet (width)
.
Length:
3w = 3(21) = 63 feet